OpenStudy (anonymous):

solve 11x=24+8x?....find value of x

OpenStudy (anonymous):

x = 8

OpenStudy (anonymous):

thankss!

OpenStudy (anonymous):

11x-8x=24 3x=24 x=24/8 x=3